A disc rotates through 10 rad in 4 seconds. If the disc start from rest, determine its angular velocity after 4s
1
Expert's answer
2021-10-03T13:23:16-0400

With the info given, find the angular acceleration:


"\alpha=\frac{2\phi}{t^2}."

The angular velocity after the 4 s is


"\omega=\alpha t=\frac{2\phi}{t}=5\text{ rad/s}."